ENCAPSULATION OF PRODAN IN A BETA-CYCLODEXTRIN ENVIRONMENT: AN EXPERIMENTAL AND THEORETICAL STUDY VIA ELECTRONIC SPECTROSCOPY AND MOLECULAR MECHANICS

ANWESHA BANERJEE, BIDISA SENGUPTA^{a}, SUDIP CHAUDHURI and \underline{PRADEEP K. SENGUPTA}, Biophysics Division, Saha Institute of Nuclear Physics, 1/AF, Bidhannagar, Kolkata 700064, India.

apresent address: Department of Applied Physics, Chalmers University of Technology, 412 96 Götenborg, Sweden